OpenStudy (anonymous):

Solve 12(1.5)^x + 4 = 13

OpenStudy (anonymous):

12 (1.5)^x = 9 (1.5)^x = 3/4 log 3/4 / log 1.5 = x

OpenStudy (mertsj):